# PUP 3D-GS: Principled Uncertainty Pruning for 3D Gaussian Splatting

### v1

@@ A.1 Scene Evaluations
LightGaussian 0.3318 0.2600 0.2828 0.4259 0.2291 0.2069 0.2616 0.3155 0.4261 0.3287 0.2039 0.3482 0.3264
Ours 0.2870 0.2304 0.2323 0.4196 0.1859 0.1569 0.2271 0.2784 0.4016 0.2981 0.1925 0.3316 0.2990
Table 9: FPS on each scene after two steps of prune-refine.
@@
Baseline (3D-GS) 50.23 130.25 123.76 106.47 60.86 98.47 115.75 87.80 86.76 167.53 128.40 90.42 120.81
LightGaussian 114.51 229.01 198.98 241.87 163.21 188.01 181.04 218.44 187.17 404.51 235.72 186.80 233.44
Ours 139.67 297.42 268.49 272.12 201.04 296.71 255.86 222.75 242.61 521.98 313.98 318.57 273.59
A.2 Additional Scene Visualizations
Figure 6: Visual comparison after two rounds of prune-refine using our method and LightGaussian’s method. Top: bicycle from the Mip-Nerf 360 dataset. Middle: stump from the Mip-Nerf 360 dataset. Bottom: playroom from the Deep Blending dataset. A larger example image of playroom can be found in Figure 1
A.3 Ablation on Patch Size
Table 10: Results of pruning the Tanks&Temples truck scene using our sensitivity score computed with $8\times 8$ and $64\times 64$ patches. Notice that, although the $8\times 8$ patches that we use in our experiments measures slightly better across all metrics, the $64\times 64$ size patches produce similar results.
Methods Tanks&Temples Truck Scene
[…]
3D-GS 25.43 0.8782 0.1476 128.40 610.19
$8\times 8$ 24.73 0.8541 0.1925 313.98 70.54
$64\times 4$ 24.63 0.8501 0.1973 312.81 70.54
A.4 Ablation on Number of Prune-Refine Steps
Table 11: Results from pruning a total of 88.44% of the Gaussians using one, two, and three steps of prune-refine over 10,000 total iterations of refinement. Our two-step method achieves the highest image fidelity.
@@ A.4 Ablation on Number of Prune-Refine Steps
Baseline (3D-GS) 25.09 0.7467 0.2442 50.23 1345.58
+ $88.44\%+10000$ 24.83 0.7177 0.2916 95.77 155.55
+ $(66\%+5000)\times 2$ 24.99 0.7270 0.2870 139.67 155.55
+ $(51.29\%+3333)\times 3$ 24.91 0.7270 0.2875 164.63 155.55
A.5 Ablation on Pruning Percentage
Table 12: Results from two steps of prune-refine with 50%, 66%, and 75% pruning and 5000 refinement iterations per-step. Our 66% pruning method balances the higher fidelity of 50% pruning with the improved FPS and file size of 75% pruning.
Methods Mip-NeRF 360 Bicycle Scene
[…]
Baseline (3D-GS) 25.09 0.7467 0.2442 50.23 1345.58
+ $(50\%+5000)\times 2$ 25.19 0.7459 0.2537 93.54 336.40
+ $(66\%+5000)\times 2$ 24.99 0.7270 0.2870 139.67 155.55
+ $(75\%+5000)\times 2$ 24.67 0.6995 0.3269 174.86 84.10
A.6 Scene Ground Truth Residuals
Figure 7: Visual comparison of our method with LightGaussian after removing $88.44\%$ of Gaussians with added ground truth L1 error visualizations. Columns "GT", "Ours", and "LightGaussian" are the same as in Figure 4. Columns "Ours L1 Error" and "LG L1 Error" are the L1 Error images of our method and LightGaussian against the ground truth images.
A.7 Scene 3D-GS Residuals
@@
Figure 8: Visual comparison after two rounds of prune-refine using our method and LightGaussian’s. with added ground truth L1 error visualizations. Columns "3D-GS", "Ours", and "LightGaussian" are the same as in Figure 4. Columns "Ours GS L1 Error" and "LG GS L1 Error" are the L1 error images of our method and LightGaussian against the original 3D-GS reconstruction of the scene.
A.8 Prune-Refining EAGLES
Table 13: Results from training EAGLES [4] on the bicycle scene and then running two steps of prune-refine. Notice that the base EAGLES model produces similar metrics to vanilla 3D-GS despite being 2.51×2.51$\times$2.51 × smaller than it. By applying PUP 3D-GS to the EAGLES model, we further reduce its size to 7.39×7.39$\times$7.39 × smaller than the vanilla 3D-GS model while achieving better image quality than LightGaussian, which attains slightly higher rendering speed.
Methods Mip-NeRF 360 Bicycle Scene
[…]
3D-GS 25.09 0.7467 0.2442 50.23 1345.58
Baseline (EAGLES) 25.07 0.7508 0.2433 58.70 535.21
EAGLES + LightGaussian 24.52 0.7018 0.3094 91.26 181.97
EAGLES + Ours 24.78 0.7212 0.2870 83.13 181.97

### v3

@@ A.4 Scene Evaluations
LightGaussian 0.3338 0.2568 0.2801 0.4258 0.2407 0.2042 0.2625 0.3132 0.4315 0.3227 0.2041 0.3383 0.3201
Ours 0.2965 0.2281 0.2297 0.4211 0.1997 0.1545 0.2278 0.2836 0.4062 0.2967 0.1916 0.3067 0.2963
Table 10: FPS on each scene after two steps of prune-refine. Results were collected with a Nvidia RTXA4000 GPU.
@@
Baseline (3D-GS) 32.32 106.65 79.70 66.48 37.60 64.46 76.31 54.04 59.08 114.45 81.26 56.79 76.79
LightGaussian 110.94 208.15 168.48 182.28 133.77 165.98 172.46 169.98 147.03 378.50 279.55 206.94 261.27
Ours 127.85 289.21 222.77 205.21 164.79 237.66 231.63 179.38 184.78 494.76 287.44 284.81 318.06
A.5 Prune-Refining EAGLES
In Table 11, we ablate our PUP 3D-GS pipeline against LightGaussian’s pipeline on an EAGLES [6] model of the Mip-NeRF 360 bicycle scene. Notice that the base EAGLES model produces similar metrics to vanilla 3D-GS despite being 2.51×2.51$\times$2.51 × smaller than it. By applying PUP 3D-GS to the EAGLES model, we further reduce its size to 25.14×25.14$\times$25.14 × smaller than the vanilla 3D-GS model while achieving better image quality and rendering speed than LightGaussian.
Table 11: Results from training EAGLES [6] on the bicycle scene and then running two steps of prune-refine with our and LightGaussian’s methods.
@@ A.5 Prune-Refining EAGLES
3D-GS 25.09 0.7467 0.2442 32.12 1345.58
Baseline (EAGLES) 25.07 0.7508 0.2433 47.82 535.21
EAGLES + LightGaussian 23.57 0.6082 0.4039 109.26 53.52
EAGLES + Ours 24.01 0.6686 0.3566 144.56 53.52
A.6 Comparison with Other Pruning Methods
Several other papers also introduce pruning techniques. Compact-3DGS[14] reports that they prune $58.7\%$ of Gaussians from the Mip-NeRF 360 bonsai scene, slightly improving PSNR from $29.87$ to $29.91$. After pruning $58.7\%$ of Gaussians with a single round of prune-refine, our method produces a much larger PSNR boost from $32.25$ to $32.55$. Mini-Splatting [5] and RadSplat [24] achieve higher rendering quality through orthogonal methods like improved densification and pretrained NeRF models. Since our pruning approach can be combined with these methods, it is unclear if a direct comparison with their results is useful. As reported in Table 2, our 10×10$\times$10 × reduction in the number of Gaussians is significantly higher than the 2.28×2.28$\times$2.28 × and 6.94×6.94$\times$6.94 × reductions reported by Compact-3DGS and Mini-Splatting.
A.7 Potential Directions for Future Research
The theory behind our Hessian approximation relies on a set of views whose L1 loss is close to zero, so we compute our spatial sensitivity score on the training views after optimization to remain as mathematically principled as possible. However, our score (1) does not rely on ground-truth data and (2) can be computed when the L1 loss is sufficiently low earlier in the 3D-GS training pipeline. These properties may provide a useful first step for several potential directions for future research.
Further optimizations and/or sufficiently powerful hardware could allow our sensitivity score to be used to prune the model during training. This is a promising research direction that could potentially lead to even higher compression and rendering speeds. Additionally, it may be possible to identify a subset of pixels, rays, or views that would produce a score that is most effective for pruning. Since this is a non-trivial, combinatorially difficult problem and our pruning score can be computed across the entire training set in seconds, we also leave this as an open research question.
Our score can be extended to other 3D Gaussian pipelines such as the Mini-Splatting and PGSR pipelines [5, 2]. Some considerations must be made, such as the effect of each Gaussian on the depth and normal maps produced by these approaches along with the rendered image. Nevertheless, the math in our main paper holds if we include the depth and normal maps as additional channels. Pruning anchor-based approaches like Scaffold-GS [17] and derived works like Octree-GS [26] is a more difficult problem. Directly pruning Gaussians with our current pipeline will not work because a fixed number of Gaussians are produced for each anchor point; pruning anchor points is ill-advised because they are generated in areas of the scene that do not have sufficient geometry. While our PUP 3D-GS approach presents a method for directly quantifying a Gaussian’s importance in the scene, determining how to map that score to pruning anchors is an open research problem.
FisherRF [9] also computes Fisher information for 3D Gaussian Splats, but uses it to perform active-view selection and post-hoc uncertainty visualization instead of pruning. Furthermore, FisherRF only approximates the diagonal of the Fisher matrix and uses the color parameters of the Gaussians, whereas our approach uses the spatial mean and scaling parameters to compute a more accurate block-wise approximation. Our sensitivity pruning score can be directly repurposed for these applications, but we leave this to future work because it is not the focus of our paper.
A.8 Additional Scene Visualizations
Figure 9 provides a visual comparison of the ground truth image against renderings from the 3D-GS model before and after pruning with our PUP 3D-GS and LightGaussian’s pipelines. Notice that our method consistently achieves higher visual fidelity and retains more salient foreground information like individual leaves and legible text.
A.9 Background Degradation
At the extreme pruning ratios used in our work, we observe some background degradation in pruned models. Intuitively, background regions that are observed from fewer viewpoints exhibit higher uncertainty than well-observed foreground regions, making them more susceptible to pruning. Given our aggressive pruning threshold of 90%, lossless compression is not expected. However, our method prioritizes preserving fine details in the foreground while pruning less important Gaussians in the background to maintain overall visual quality. We do not consider this a failure case because retaining foreground details over background details is often preferable.
In contrast, LightGaussian exhibits the opposite tendency, often favoring background preservation over foreground fidelity. Figure 10 illustrates this difference by comparing L1 residuals against renderings from the base 3D-GS model and highlighting instances where PUP 3D-GS degrades the background more than LightGaussian. A potential strategy for mitigating this trade-off is to reweight the background Gaussian scores using masking, incorporating a tunable parameter to control the balance between foreground and background degradation. We leave this exploration to future work.
A.10 Scene Residuals
Figures 11 and 12 provide visual comparisons of the L1 residual of renderings from the 3D-GS model before and after pruning $90\%$ of Gaussians with our PUP 3D-GS and LightGaussian’s pipelines. Figure 11 compares the L1 residuals with respect to the ground truth image, while Figure 12 compares them with respect to renderings from the base 3D-GS model. In both cases, our PUP 3D-GS pipeline produces less L1 error and retains more salient foreground information than LightGaussian’s.
Figure 9: Visual comparison after two rounds of prune-refine using our method and LightGaussian’s method. Top: bicycle from the Mip-Nerf 360 dataset. Middle: stump from the Mip-Nerf 360 dataset. Bottom: playroom from the Deep Blending dataset. A larger example image of playroom can be found in Figure 1.
Figure 10: Cases where our method produces more background L1 error against renderings from the base 3D-GS model than LightGaussian. Columns "3D-GS", "Ours", and "LightGaussian" are the same as in Figure 4. Columns "Ours GS L1 Error" and "LG GS L1 Error" are the L1 error images of our method and LightGaussian against the original 3D-GS reconstruction of the scene. Our method prioritizes retaining visual quality and fine details in the foreground over preserving less important information in the background.
@@ A.10 Scene Residuals
Figure 11: Visual comparison after two rounds of prune-refine using our method and LightGaussian’s with additional L1 error visualizations against the ground truth images. Columns "GT", "Ours", and "LightGaussian" are the same as in Figure 4. Columns "Ours L1 Error" and "LG L1 Error" are the L1 Error images of our method and LightGaussian against the ground truth images. Our method produces lower error than LightGaussian in all examples.
Figure 12: Visual comparison after two rounds of prune-refine using our method and LightGaussian’s with additional L1 error visualizations against renderings from the base 3D-GS model. Columns "3D-GS", "Ours", and "LightGaussian" are the same as in Figure 4. Columns "Ours GS L1 Error" and "LG GS L1 Error" are the L1 error images of our method and LightGaussian against the original 3D-GS reconstruction of the scene. Our method produces lower error than LightGaussian in all examples.